# Can we Evaluate RAGs with Synthetic Data?

### v1

@@ Abstract
We investigate whether synthetic question-answer (QA) data generated by large language models (LLMs) can serve as an effective proxy for human-labeled benchmarks when such data is unavailable. We assess the reliability of synthetic benchmarks across two experiments: one varying retriever parameters while keeping the generator fixed, and another varying the generator with fixed retriever parameters. Across four datasets, of which two open-domain and two proprietary, we find that synthetic benchmarks reliably rank the RAGs varying in terms of retriever configuration, aligning well with human-labeled benchmark baselines. However, they fail to produce consistent RAG rankings when comparing generator architectures. The breakdown possibly arises from a combination of task mismatch between the synthetic and human benchmarks, and stylistic bias favoring certain generators.
Keywords:
Retrieval-Augmented Generation Synthetic Benchmarking Question-Answering Evaluation Large Language Models
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has been proposed as a method for mitigating inaccurate and contextually inappropriate responses in domain-specific applications of Large Language Models (LLMs) [11, 20, 15]. By incorporating factual information from external documents into the prompt, RAG models enhance explainability and domain specificity whilst reducing hallucinations [12, 13, 10, 9]. Commercial products such as Pega Knowledge Buddy (KB) have leveraged this technology to facilitate a simple way of creating domain-specific QA systems suited to a wide range of internal and external use cases, including human resources management, customer service, sales, and due diligence. Since the relationship between RAG configuration and its downstream performance is complex, RAGs require rigorous benchmarking, especially in regulatory, cost- and error-sensitive environments [1, 5].
@@
Instead, selected domain-specific research has opted for synthetic QA datasets [8, 14, 7, 23, 26, 18, 6]. Generating synthetic QA pairs based on domain knowledge enables the construction and reconstruction of an up-to-date domain-specific dataset with pseudo-ground-truth answers. Although synthetic benchmarks pose an attractive solution for the automated evaluation of domain-specific RAG models, their reliability has not been systematically assessed in the context of design choices made with real benchmark data, raising concerns about the ecological validity of the method. To address this gap, we structure our study around the following two research questions:
  1. 1.
    How well do synthetic QA benchmarks align with human-annotated benchmarks in ranking question-answering RAG systems on quality?
  2. 2.
    Can synthetic benchmarks serve as an effective method for automated evaluation of domain-specific RAG models?
@@ 3 Methods
RAG architectures operate by chunking corpus documents into overlapping fixed-size segments and storing these in a knowledge base. At query time, a dense retriever identifies the top-$k$ semantically relevant chunks, subject to a semantic similarity threshold, which are then injected into a prompt template that is fed to the LLM for answer generation.
@@ 4 Experiments
This section describes the experimental conditions under which we test the validity of synthetic benchmarks for RAG systems. We detail the default individual model configuration and evaluation, datasets used, the synthetic data generation process, model variations evaluated, and our benchmark evaluation procedure. [^1] Code is available at https://github.com/JonasElburgUVA/Can-we-evaluate-RAGs-with-synthetic-data/blob/master.
@@ 4.2 RAG scoring
The resulting RAG systems are evaluated in terms of their answer accuracy compared to a reference answer, using the Ragas framework [8]. We assess answer accuracy using the key-word based ROUGE-L F1, BLEU, and Levenshtein string similarity, alongside semantic similarity measured with OpenAI text-embedding-3-small. Additionally, three LLM-based evaluation metrics from the Ragas framework are implemented: Answer relevance, Faithfulness, and Context Precision (LLM-based MAP@k), all using GPT-4o-mini and the prompts provided by Ragas to assess different aspects of the given answer [8]. More details on these metrics can be found in the supplementary materials.
@@ 4.4 Synthetic QA generation
To generate synthetic QA pairs based on both datasets, we adapt prompts described by LlamaIndex [24], requesting questions and answers that must be answerable given a document from our corpora (Supplementary materials Listing 2). The LLM is GPT-4o version 2024-11-20 with temperature set to 0.7 and top-$p$ set to 0.95 (Supplementary materials Listing 2). We exclude self-validating techniques such as round-trip consistency to obtain an unbiased baseline performance of naive synthetic data generation.
We sample 100 synthetic QA pairs per dataset using a default of 20 documents, meaning we obtain five questions per randomly sampled document. For the Launchpad dataset this is adjusted to two questions per document due to a lack of lengthy documents in the corpus (Table 2). For the Sales dataset, which has a similarly low average document length but far more total available documents, we instead employ a minimum document length of 500 words.
@@
Document Length (words) 4319 $\pm$ 1678 3522 $\pm$ 4284 488 $\pm$ 405 645 $\pm$ 1146
@@ 4.6 Evaluation
Given a set of questions, retrieved contexts, generated answers and reference answers we calculate the evaluation metrics described in section 4.2 for each model. The evaluation metrics are used to construct rankings of the RAG systems according to the human and synthetic benchmarks. We then calculate the Kendall rank correlation between the two rankings for each metric of interest, providing a statistic reflecting how well the synthetic benchmark matches the human benchmark under each metric. A perfectly consistent ranking produces $\tau=1.0$, whereas a completely inverted ranking produces $\tau=-1.0$. The results are not evaluated for statistical significance due to the limited number of models evaluated.
@@ 5.2 Experiment A - Retrieval
We observe moderate to high ranking consistency between human and synthetic benchmarks across all datasets and evaluation metrics (Table 5). On average, the SQuAD benchmarks shows the lowest average rank correlation (mean Kendall’s $\tau=0.44$), while the ASQA benchmarks exhibit the highest ($\tau=0.75$). Figure 1 indicates that ASQA has a lower task mismatch between benchmarks, while SQuAD and Sales may have suffered from a higher task mismatch than ASQA and Launchpad. These results suggest that ranking agreement between synthetic and human benchmarks is to an extent dataset-dependent.
@@
Considering the different metrics used, we find the highest ranking consistency using the keyword-based and precision-oriented BLEU metric ($\tau=0.84$). ROUGE (keyword-based, recall-oriented) and Semantic Similarity (embedding-based) also produce strong alignment ($\tau=0.67$), suggesting that both surface-level and semantic supervised measures of correctness can yield consistent rankings.
The unsupervised Faithfulness and Answer Relevance metrics produce consistent rankings across datasets, and prefer higher context-windows, similar to more standard supervised metrics such as semantic similarity. In contrast, Context Precision reveals notable discrepancies between human and synthetic benchmarks ($\tau=0.08$). In synthetic datasets, Context Precision scores are high across all retrieval settings, indicating the most relevant chunk is often ranked highest by the semantic search and is sufficient for answering the question (Figure 3). In contrast, most human benchmarks show a clear performance gap: the low-context model consistently underperforms, indicating that single top-ranked documents are often insufficient for answering questions. This again suggests that the synthetic benchmark may underestimate the retrieval challenge posed by natural user queries.
Table 5: Kendall’s $\tau$ rank correlation coefficients between human and synthetic benchmarks across four datasets (rows = metrics, columns = datasets). SQuAD showed the lowest consistency across metrics, while Sales yielded the highest. BLEU and Semantic Similarity achieved the most consistent rankings overall.
SQuAD ASQA Launchpad Sales Mean
BLEU 0.67 1.00 1.00 0.67 0.84
ROUGE 0.67 0.67 0.67 0.67 0.67
Levenshtein Similarity 0.33 0.67 0.67 0.33 0.50
Semantic Similarity 0.33 1.00 0.67 0.67 0.67
False negative rate 0.18 0.67 0.91 0.80 0.64
Faithfulness 0.67 0.67 0.67 1.00 0.75
Context Precision 0.33 0.67 0.00 -0.67 0.08
Answer Relevance 0.33 0.67 0.67 1.00 0.67
Mean 0.44 0.75 0.66 0.56
@@ 5.3 Experiment B - LLMs
Experiment B evaluates the consistency of model rankings across five large language models (LLMs) as part of RAG systems using both human and synthetic evaluation benchmarks. Retrieval settings are held constant to isolate the effect of the generator model choice.
Overall, rankings differ substantially between the synthetic and human benchmarks across datasets and evaluation metrics (Table 6). Several metric–dataset combinations yield low or even negative Kendall’s $\tau$ values, indicating little agreement and, in some cases, invert model preferences. These inconsistencies are observed across both supervised metrics and unsupervised LLM-based metrics such as Faithfulness and Answer Relevance, suggesting that the ranking differences are not solely due to variation in reference answers.
Table 6: Kendall’s $\tau$ rank correlation coefficients between human and synthetic benchmarks across four datasets (rows = metrics, columns = datasets).
SQuAD ASQA Launchpad Sales Mean
BLEU 0.00 -0.80 0.40 -1.00 -0.35
ROUGE 0.60 -0.60 -0.20 -1.00 -0.30
Levenshtein Similarity 0.60 -0.60 -0.20 -1.00 -0.30
Semantic Similarity 0.40 -0.20 0.20 0.00 0.10
False Negative Rate 0.33 0.80 -0.35 1.00 0.45
Faithfulness 1.00 0.60 -0.20 0.00 0.35
Context Precision -0.60 0.80 0.00 -0.67 -0.12
Answer Relevance 0.00 0.40 -0.20 0.00 0.05
Mean 0.29 0.05 -0.07 -0.33
@@ 6.1 Reflections on results
Experiment A demonstrates moderate to high ranking consistency between human and synthetic benchmarks across datasets and evaluation metrics, with the exception of the Context Precision metric. In both benchmark types, larger context windows generally correlated with higher scores, suggesting that additional input, even if only partially relevant, positively impacts model performance. The ranking consistency is highest for ASQA and Launchpad data, which likely contained less task mismatch originating from question and answer lengths than SQuAD and Sales (Figure 1). These results suggest that synthetic benchmarks with minimized task mismatches can be used when optimizing retrieval parameters.
@@ 6.2 Known limitations
Our assessed synthetic benchmark is limited to one QA generation method, without considering more complex synthetic QA pairs such as multi-hop or adversarial questions [2]. Furthermore, the prompt used to synthesize QA pairs is relatively simplistic and would likely benefit from custom few-shot examples, as was done in previous work [25, 13, 21]. Additionally, we did not use self-validation methods such as round-trip consistency to filter the synthetic data, which may increase the reliability of the benchmark [3, 21]. With additional calibration these factors could mitigate task mismatch and influence ranking consistency.

### v2

@@ Abstract
We investigate whether synthetic question-answer (QA) data generated by large language models (LLMs) can serve as an effective proxy for human-labeled benchmarks when the latter is unavailable. We assess the reliability of synthetic benchmarks across two experiments: one varying retriever parameters while keeping the generator fixed, and another varying the generator with fixed retriever parameters. Across four datasets, of which two open-domain and two proprietary, we find that synthetic benchmarks reliably rank the RAGs varying in terms of retriever configuration, aligning well with human-labeled benchmark baselines. However, they do not consistently produce reliable RAG rankings when comparing generator architectures. The breakdown possibly arises from a combination of task mismatch between the synthetic and human benchmarks, and stylistic bias favoring certain generators.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has been proposed as a method for mitigating inaccurate and contextually inappropriate responses in domain-specific applications of Large Language Models (LLMs) [11, 20, 15]. By incorporating factual information from external documents into the prompt, RAG models enhance explainability and domain specificity whilst reducing hallucinations [12, 13, 10, 9]. Commercial products such as Pega Knowledge Buddy (KB), on which this study is based, leverage this technology to facilitate a simple way of creating domain-specific question answering (QA) systems suited to a wide range of internal and external use cases, from human resource management, customer service, sales, to customer due diligence. Since the relationship between RAG configuration and its downstream performance is complex, RAGs require rigorous benchmarking, especially in regulatory, cost- and error-sensitive environments [1, 5].
@@
Instead, selected domain-specific research has opted for synthetic QA datasets [8, 14, 7, 23, 26, 18, 6]. Generating synthetic QA pairs based on domain knowledge enables the construction and reconstruction of an up-to-date domain-specific dataset with pseudo-ground-truth answers. Although synthetic benchmarks pose an attractive solution for the automated evaluation of domain-specific RAG models, their reliability has not been systematically assessed in the context of design choices made with real benchmark data, raising concerns about the ecological validity of the method. To address this gap, we structure our study around the following two research questions: how well do synthetic QA benchmarks align with human-annotated benchmarks in ranking question-answering RAG systems on quality, and can synthetic benchmarks serve as an effective method for automated evaluation of domain-specific RAG models.
@@ 3 Methods
RAG architectures operate by chunking corpus documents into overlapping fixed-size segments and storing these in a knowledge base. At query time, a dense retriever identifies the top-$k$ semantically relevant chunks, subject to a semantic similarity threshold, which are then injected into a prompt template that is fed to the LLM for answer generation [12].
@@ 4 Experiments
This section describes the experimental conditions under which we test the validity of synthetic benchmarks for RAG systems. We detail the default individual model configuration and evaluation, datasets used, the synthetic data generation process, model variations evaluated, and our benchmark evaluation procedure. [^1] Code and supplementary materials are available at https://github.com/JonasElburgUVA/Can-we-evaluate-RAGs-with-synthetic-data/blob/master.
@@ 4.2 RAG scoring
The resulting RAG systems are evaluated in terms of their answer accuracy compared to a reference answer using the Ragas framework [8]. We assess answer accuracy using the key-word based ROUGE-L F1, BLEU, and Levenshtein string similarity, alongside semantic similarity measured with OpenAI text-embedding-3-small. Additionally, three LLM-based evaluation metrics from the Ragas framework are implemented: Answer relevance, Faithfulness, and Context Precision (LLM-based MAP@k), all using GPT-4o-mini and the prompts provided by Ragas to assess different aspects of the given answer [8]. More details on these metrics can be found in the supplementary materials.
@@ 4.4 Synthetic QA generation
To generate synthetic QA pairs based on both datasets, we adapt prompts described by LlamaIndex [24], requesting questions and answers that must be answerable given a document from our corpora (Supplementary materials Listing 2). The LLM is GPT-4o version 2024-11-20 with temperature set to 0.7 and top-$p$ set to 0.95. We exclude self-validating techniques such as round-trip consistency to obtain an unbiased baseline performance of naive synthetic data generation.
We sample 100 synthetic QA pairs per dataset using a default of 20 randomly sampled documents, meaning we obtain five questions per document. For the Launchpad dataset this is adjusted to two questions per document and 50 sampled documents due to a lack of lengthy documents in the corpus (Table 2). For the Sales dataset, which has a similarly low average document length but far more total available documents, we instead employ a minimum document length of 500 words.
@@
Words per Document ($\mu\pm\sigma$) 4319 $\pm$ 1678 3522 $\pm$ 4284 488 $\pm$ 405 645 $\pm$ 1146
@@ 4.6 Evaluation
Given a set of questions, retrieved contexts, generated answers and reference answers we calculate the evaluation metrics described in section 4.2 for each model. The evaluation metrics are used to construct rankings of the RAG systems according to the human and synthetic benchmarks. We then calculate the Kendall rank correlation between the two rankings for each metric of interest, providing a statistic reflecting how well the synthetic benchmark matches the human benchmark under each metric. A perfectly consistent ranking produces $\tau=1.0$, whereas a completely inverted ranking produces $\tau=-1.0$. The statistic is not evaluated for statistical significance due to the limited number of models evaluated.
@@ 5.2 Experiment A - Retrieval
We observe moderate to high ranking consistency between human and synthetic benchmarks across all datasets and evaluation metrics, with the SQuAD dataset resulting in slightly lower scores than other datasets (Figure 2). The difference between synthetic and human QA length observed in Figure 1 indicates that SQuAD and Sales may have suffered from a higher format mismatch than ASQA and Launchpad. These results suggest that ranking agreement between synthetic and human benchmarks is to an extent dataset-dependent.
@@
Considering the different metrics used, we find the highest average ranking consistency using the keyword-based and precision-oriented BLEU metric ($\tau_{\mu}=0.84$). ROUGE (keyword-based, recall-oriented) and Semantic Similarity (embedding-based) also produce strong alignment ($\tau_{\mu}=0.67$), suggesting that both surface-level and semantic supervised measures of correctness can yield consistent rankings.
The unsupervised Faithfulness and Answer Relevance metrics produce consistent rankings across datasets, and prefer higher context-windows, similar to more standard supervised metrics such as semantic similarity. In contrast, Context Precision reveals notable discrepancies between human and synthetic benchmarks ($\tau_{\mu}=0.08$). In synthetic datasets, Context Precision scores are high across all retrieval settings, indicating the most relevant chunk is often ranked highest by the semantic search and is sufficient for answering the question (Figure 4). In contrast, most human benchmarks show a clear performance gap: the low-context model consistently underperforms, indicating that single top-ranked documents are often insufficient for answering questions. This again suggests that the synthetic benchmark may underestimate the retrieval challenge posed by natural user queries.
Refer to caption Refer to caption
Figure 2: Kendall’s $\tau$ rank correlation coefficients between human and synthetic benchmarks across four datasets and two experiments. Rank correlation is generally higher in the retriever experiment than the generator experiment. Furthermore, consistency is to an extent dataset- and metric-dependent.
@@ 5.3 Experiment B - LLMs
Experiment B evaluates the consistency of model rankings across five large language models (LLMs) as part of RAG systems using both human and synthetic evaluation benchmarks. Retrieval settings are held constant to isolate the effect of the generator model choice (Figure 5 and 6).
Overall, rankings differ substantially between the synthetic and human benchmarks across datasets and evaluation metrics (Figure 2). Several metric–dataset combinations yield low or even negative Kendall’s $\tau$ values, indicating little agreement and, in some cases, invert model preferences. These inconsistencies are observed across both supervised metrics and unsupervised LLM-based metrics such as Faithfulness and Answer Relevance, suggesting that the ranking differences are not solely due to variation in reference answers.
Figure 5: Supervised metrics in the LLM choice experiment. Error bars represent the variance over the 94-100 datapoints. The Llama model is missing from the Sales experiment since it was removed from the KB platform since the start of the research project.
Figure 6: Unsupervised metrics in the LLM choice experiment. Error bars represent the variance over the 94-100 datapoints. The Llama model is missing from the Sales experiment since it was removed from the KB platform since the start of the research project.
@@ 6.1 Reflections on results
Experiment A demonstrates moderate to high ranking consistency between human and synthetic benchmarks across datasets and evaluation metrics, with the exception of Context Precision. In both benchmark types, larger context windows generally correlates with higher scores, suggesting that additional input, even if only partially relevant, positively impacts model performance. The ranking consistency is highest for ASQA and Launchpad data, which likely contains less task mismatch originating from question and answer lengths than SQuAD and Sales (Figure 1). These results suggest that synthetic benchmarks with minimized task mismatches can be used when optimizing retrieval parameters.
@@ 6.2 Known limitations
We purposely used only a single, simple synthetic QA generation method, without considering more complex synthetic QA pairs such as multi-hop or adversarial questions [2]. Furthermore, the prompt used to synthesize QA pairs is relatively simplistic and would likely benefit from custom few-shot examples, as was done in previous work [25, 13, 21]. Additionally, we did not use self-validation methods such as round-trip consistency to filter the synthetic data, which may increase the reliability of the benchmark [3, 21]. With additional calibration these factors could mitigate task mismatch and influence ranking consistency.